Hankins, Inc., is considering a project that will result in initial aftertax cash savings of $5.7 million at the end of the first year, and these savings will grow at a rate of 3 percent per year indefinitely. The firm has a target debt-equity ratio of .56, a cost of equity of 13.1 percent, and an aftertax cost of debt of 5 percent. The cost-saving proposal is somewhat riskier than the usual project the firm undertakes; management uses the subjective approach and applies an adjustment factor of +2 percent to the cost of capital for such risky projects. a. Calculate the required return for the project. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What is the maximum cost the company would be willing to pay for this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Given, Debt-Equity Ratio= 0.56.

Therefore, proportion of debt (Wd)=0.56 and Proportion of Equity (We)= 1-0.56= 0.44

Also given, Cost of equity (Re) = 13.1%, After-tax cost of debt (Rd)=5%

Weighted average cost of capital WACC= Wd*Rd +We*Re

                                                                          = 0.56*5% + 0.44*13.1% = 2.8% + 5.764% = 8.564%

Given that adjustment factor of +2 percent is applied on the cost of capital considering the risk involved. Therefore, required rate of return= WACC + 2% = 8.564% + 2% = 10.56% (rounded to 2 decimals)

Maximum cost of the project that the company is willing to pay is the present value of future cash flows, discounted at the required rate of return.

Given,

First year after-tax cash savings=$5.7 Million

Growth rate of yearly cash savings= 3% (Infinite)

Present value of perpetual annuity when there is constant growth is ascertained using the formula

PV = P/(r-g)

Where PV= Present Value, P= Initial cash flow (given as $5.7 Million), r= Discount rate (or the rate of return) arrived at 10.564% and g= constant growth rate given as 3%.

Applying the values in the formula, Present Value= $5.7 Million/(10.564%-3%)

                               = 5.7/0.10564-0.03) = 5.7/7.564 = $75.356954 Million (Rounded to $75.36 Million)

Therefore, the Maximum cost the company is willing to pay for this project is $75.36 Million